OpenStudy (anonymous):

the question is: find each expression containing the principal square root. (then) sqrt 10+15?

OpenStudy (lgbasallote):

\[\sqrt{10} + 15\] or \[\sqrt{10+15}\]

OpenStudy (anonymous):

the latter

OpenStudy (anonymous):

2nd one

OpenStudy (lgbasallote):

\[\sqrt{10+15} = \sqrt{25} = \pm 5\] since you're looking for principal square root it will just be POSITIVE 5

OpenStudy (anonymous):

oh okay ty, I was way off...thank you so very much

OpenStudy (lgbasallote):

welcome :D